# We can catalyse it for you thermally: Catalytic transformations for thermal operations

### v1

@@ IV Quantifying catalytic advantages
[…]
@@
V Entanglement generation under thermal operations with catalyst
@@
Recently, the necessary and sufficient conditions for producing bipartite qubit entanglement from an initially separable states via thermal operations were addressed in Ref. [48]. States that cannot get entangled under thermal operations are said to belong to the thermally non-entanglable set, mathematically denoted by $\mathbb{TN}$. Conversely, states that can get entangled under thermal operations lie in the entanglable set $\mathbb{TE}$. By utilising the results on the catalysable future region $\mathcal{C}_{+}$, we may now answer the natural question whether borrowing catalyst allow more states to get entangled. Defining the catalytically non-entanglable set $\mathbb{CN}$ as the set of states that cannot be entangled even with the aid of catalysis, the question can be reframed – is the set $\mathbb{CN}$ smaller than $\mathbb{TN}$, or in terms of volumes, $\mathcal{V}(\mathbb{CN})/\mathcal{V}(\mathbb{TN})\overset{?}{<}1$ for a given $\beta$?
@@
Using the introduced numerical approximations we derived the dependence of the volumes $\mathcal{V}(\mathbb{TN})$ and $\mathcal{V}(\mathbb{CN})$ for both thermall and catalytically non-entanglable sets and their ratio as a function of $\beta$, demonstrated in Fig 7.
@@
VI Summary and Discussion
@@
In this paper, we explore the fundamental limits on state achievability with the aid of a strict catalyst under a thermal process. We propose a novel approach to address this question by constructing the sets of states that can be achieved from a given energy-incoherent state under catalytic thermal operations, as well as the sets of states that can, through catalytic thermal operations, achieve the initial state. These regions naturally highlight the advantages of using a strict catalyst in general thermodynamic processes, whose only assumption is energy conservation. Our construction enables us to characterise the catalysable past and future for main systems of arbitrary dimensions. This analytical approach is based solely on thermomajorisation relations. Furthermore, we established general bounds on the catalyst dimension for thermal operations and demonstrated that catalysis is effective under thermal operations even for systems of dimensionality as small as $d=3$. The catalytic advantages were quantified by the volume of the catalysable future, with a detailed discussion on its behaviour as a function of temperature. Finally, we applied our findings to a practical question: the generation of entanglement under thermal operations. We showed that thermal processes, which were once incapable of generating entanglement, allows its generation when a strict catalyst is employed.
@@
To address this problem, we overcome several challenges. Firstly, there is no efficient method for characterizing the set of allowed state transformations under catalytic thermal operations. A comprehensive analysis for inexact catalysis (where the catalyst is allowed to be returned with some minor error) was performed in [25]; while, in Ref. [30], the authors solved the characterisation problem for three-dimensional systems and for a subset of initial states of generic dimensions in the context of catalysis in elementary thermal operations. Secondly, keeping track of the the set of achievable state after “decoupling” the catalyst without specifying the precise form of the catalyst state required by the transformation is a highly difficult mathematical problem. Approaching from the perspective of lattice theory, this problem can be framed by asking (i) how to characterise the higher-dimensional lattice for composite systems (main system plus catalyst) and (ii) how to project back to the native lattice of the main system while preserving the catalytic effects in the native lattice of the initial state. For small-dimensional systems (such as $d=3$), the entire convex polytope $\boldsymbol{p}\otimes\boldsymbol{c}$ can be constructed, and intersections between such a polytope and plane equations enforcing the catalyst recovery condition can be numerically identified. However, this method does not scale well and depends heavily on the initial state of the system and the catalyst. In contrast, our approach requires only constructing a set of quasi-probability vectors or determining the extremal points through thermomajorisation relations.
@@
There are many possibilities for extending our results. For instance, they can be leveraged to study catalysis within the framework of elementary thermal operations. Note that the characterisation of the catalysable past and future relies entirely on the concept of tangent vectors, which tightly thermomajorise $\boldsymbol{p}$. On the other hand, the extreme points of set of states achievable under elementary thermal operations are known to be $\beta$-swaps that tightly thermomajorise the initial distribution. Consequently, one could explore how to adapt the technical aspects of our findings to this specific context. In a similar spirit, we could relax the condition requiring the catalyst to be returned unperturbed and allow for some small error. A natural quantifier for this error is the trace distance, and the problem could be reframed by asking for the set of states achievable with a given catalyst up to this specified amount of error [25].
@@ Acknowledgements.
We thank Karol Życzkowski for asking a stimulating question at the beginning of the authors’ PhD, which led to this work, and for his fruitful comments on the first version of this manuscript. We also thank Kamil Korzekwa for valuable discussions that significantly contributed to the execution of this project, as well as for his insightful comments during the final stages of preparation of this manuscript. JCz acknowledges financial support by NCN PRELUDIUM BIS no. DEC-2019/35/O/ST2/01049 and ID.UJ Research Support minigrant. AOJ acknowledges financial support from VILLUM FONDEN through a research grant (40864).
@@ B.2 Embedding map approach
[…]

### v2

@@ III.3 Quantifying catalytic advantages
$\boldsymbol{g}=\frac{1}{Z}_{\boldsymbol{g}}\left(e^{-\beta E_{1}},...,e^{-\beta E_{d-1}},0\right)\quad\text{where}\quad Z_{\boldsymbol{g}}=\sum_{i=1}^{d-1}e^{-\beta E_{i}},$ (44)
@@
IV Applications
IV.1 Entanglement generation under thermal operations with catalyst
@@
Recently, the necessary and sufficient conditions for producing bipartite qubit entanglement from an initially separable states via thermal operations were addressed in Ref. [46]. States that cannot get entangled under thermal operations are said to belong to the thermally non-entanglable set, mathematically denoted by $\mathbb{TN}$. Conversely, states that can get entangled under thermal operations lie in the entanglable set $\mathbb{TE}$. By using the results on the catalysable future region $\mathcal{C}_{+}$, we may now answer the natural question whether borrowing catalyst allow more states to get entangled. Defining the catalytically non-entanglable set $\mathbb{CN}$ as the set of states that cannot be entangled even with the aid of catalysis, the question can be reframed – is the set $\mathbb{CN}$ smaller than $\mathbb{TN}$, or in terms of volumes, $\mathcal{V}(\mathbb{CN})/\mathcal{V}(\mathbb{TN})\overset{?}{<}1$ for a given $\beta$?
@@
Using the introduced numerical approximations we derived the dependence of the volumes $\mathcal{V}(\mathbb{TN})$ and $\mathcal{V}(\mathbb{CN})$ for both thermal and catalytically non-entanglable sets and their ratio as a function of $\beta$, demonstrated in Fig 7.
@@
IV.2 Optimal cooling with a catalyst
As a second application of our findings, we explore the task of cooling a three-level system with the aid of a catalyst. Our setup comprises a three-level system, a thermal environment, and a catalyst. We assume that the composite system forms a closed system and evolves unitarily under an energy-preserving interaction (as discussed in Sec. II). Consequently, the energy exchange in the process is account for as heat, which can be determined by calculating the energy difference between the initial and final states after the interaction. Thus, our figure of merit is defined as follows
[…]
where $\rho$ and $\sigma$ are energy-incoherent states. In other words, we ask for the optimal thermal operation which cools down the system $Q_{c}(\rho)<0$, while heating up the environment. As discussed in Sec. II, the set of achievable states under thermal operation is convex and one can prove that the optimal cooling (when it happens) occurs when the initial state is mapped to one of the extreme points of its future thermal cone. Now one can ask, what is the optimal cooling that can be achieved when we allow a strict catalyst? This question can be answered by employing Theorem 2 and evaluating the extreme point of the catalysable set. Since answering such a question in a general manner is highly non-trivial due to the state-dependent character of the problem, we can assume that we have access to the initial state preparation of the system and use our results to find the optimal cooling.
Consider, for instance, an initial state denoted by $\boldsymbol{p}=(0.1,0.2,0.7)$, characterised by a Hamiltonian with an equidistant spectrum: $E_{0}=0$, $E_{1}=1$, and $E_{2}=2$. Additionally, let us assume that the system can interact with a thermal environment prepared at an inverse temperature of $\beta=0.2$. This state has $\beta$-ordering $\boldsymbol{\pi}_{\boldsymbol{p}}=(3,2,1)$. To effectively cool the system, the optimal strategy is to send it to the extreme point of its future thermal cone with $\beta$-ordering $\boldsymbol{\pi}^{\star}=(1,2,3)$. Consequently, the initial state $\boldsymbol{p}$ transforms to $\boldsymbol{p}^{\boldsymbol{\pi}^{\star}}\approx(0.78,0.15,0.07)$. This transformation results in a heat exchange of approximately $Q_{c}\approx-1.3135$, effectively cooling the system down.
Allowing a strict catalyst to aid the process, as given in Theorem 2, we observe a shift in the extreme point with $\beta$-ordering $\boldsymbol{\pi}^{\star}=(1,2,3)$. This shift allows for a more “substantial cooling”. According to Corollary 3, our results suggest that the initial state $\boldsymbol{p}$ could now be transformed to $\boldsymbol{p}^{\star}\approx(0.85,0.08,0.07)$ (see Fig. 8). In this process, the heat exchange is approximately $Q^{\star}_{c}\approx-1.38$, surpassing $|Q_{c}|$ in magnitude. This simple example illustrates the fundamental limits for cooling a quantum system under thermal operations.
Figure 8: Catalytic cooling. For a three-level system with a population given by $\boldsymbol{p}=(0.1,0.2,0.7)$, represented by a red dot, and energy spectrum $E_{1}=0,E_{2}=1,E_{3}=2$ for $\beta=0$, the optimal cooling (without a catalyst) consists of mapping $\boldsymbol{p}$ to the extreme point of its future thermal cone with $\beta$-ordering (1, 2, 3) (dark blue point). When a catalyst is allowed, its future thermal cone expands, and the optimal cooling is achieved by sending $\boldsymbol{p}$ to the extreme point of the catalysable set with the same $\beta$-ordering (light blue dot). The arrow from the dark blue point simply indicates that a catalyst allows for bringing the initial state closer to the ground state compared to the case without a catalyst.
A more general statement can be made by considering a $d$-dimensional system described by an equidistant Hamiltonian $H=\sum_{n=0}^{d-1}n\outerproduct{n+1}{n+1}$ |, with $n\in\quantity{0,d-1}$, in contact with a thermal bath at an inverse temperature $\beta$. Assuming that the system is prepared in a thermal state at some hotter inverse temperature $\beta_{\text{h}}<\beta$, its $\beta$-ordering is $(d,d-1,\ldots,1)$ and its slope vector has entries given by $s(\boldsymbol{p})i=\exp[\quantity(\beta-\beta_{\text{h}})E_{d-i}]/Z_{\text{h}}Z$, where $Z_{\text{h}}=\sum_{i=0}^{d-1}e^{-\beta_{\text{h}}E_{i}}$. The optimal cooling is generically achieved by transforming the state $\boldsymbol{p}$ into a state $\boldsymbol{q}$ with $\beta$-order $(1,2,\ldots,d)$. Observing that for each $j\in\{0,d-1\}$, there exists an $m_{j}\in\{0,d-1\}$ such that the inequality
$\sum_{i=1}^{m_{j}}\gamma_{d-i+1}\leq\sum_{i=1}^{j}\gamma_{i}\leq\sum_{i=1}^{m_{j}+1}\gamma_{d-i+1}$ (49)
holds, we can state that $\sum_{i=1}^{m_{1}}p_{d-i+1}\leq q_{1}\leq\sum_{i=1}^{m_{1}+1}p_{d-i+1}$ is valid for the population of $\boldsymbol{q}$ corresponding to the ground state. On the other hand, we can consider the catalytic transformation $\boldsymbol{p}\otimes\boldsymbol{c}\rightarrow\boldsymbol{q}_{c}\otimes\boldsymbol{c}$, where $\boldsymbol{q}_{c}$ represents the optimal target state achieved with the catalyst. Taking into account Theorem 2, we find that
$(\boldsymbol{q}_{c})_{1}\leq s(\boldsymbol{p})_{1}\gamma_{1}=\frac{1}{Z_{\text{h}}}e^{\quantity(\beta-\beta_{\text{h}})E_{d}-\beta E_{1}}.$ (50)
Given our assumption that the system Hamiltonian is described by an equidistant spectrum, we ensure that $(\boldsymbol{q}_{c})_{1}\geq q_{1}$ when the following inequality holds:
$\frac{1-e^{-\beta d}}{1-e^{-\beta}}\frac{1-e^{-(m_{j}+1)\beta_{\text{h}}}}{1-e^{-\beta_{\text{h}}}}\leq e^{(\beta-\beta_{\text{h}})(d-1)}.$ (51)
While there is no closed-form solution to the above inequality, one can still numerically evaluate critical temperatures $\beta^{\downarrow}_{\text{h}}$, for which the catalytic advantage is guaranteed. Conversely, if
$e^{(\beta-\beta_{\text{h}})(d-1)}\leq\frac{1-e^{-\beta d}}{1-e^{-\beta}}\frac{1-e^{-m_{j}\beta_{\text{h}}}}{1-e^{-\beta_{\text{h}}}},$ (52)
we obtain a no-go result indicating that no catalytic advantage for cooling starting from a hot Gibbs state can be achieved above a certain inverse temperature $\beta^{\uparrow}_{\text{h}}$.
For very high temperatures, i.e., when $\beta_{\text{h}}<\beta\ll 1$, we can expand both $\beta$ and $\beta_{\text{h}}$ in Eqs. (51)-(52) to linear order to obtain approximate critical temperatures as
$\beta^{\downarrow}_{\text{h}}$ $=\frac{d\quantity[3\beta(d-1)-2](m_{j}+1)+2}{2d-m_{j}-2},$ (53)
$\beta^{\uparrow}_{\text{h}}$ $=\frac{d\quantity[3\beta(d-1)-2]m_{j}+2}{2d-m_{j}-1}.$
Consequently, catalytic advantages in cooling the initial system are guaranteed for $\beta_{h}\leq\beta^{\downarrow}{\text{h}}$. Conversely, for $\beta_{h}\leq\beta^{\uparrow}{\text{h}}$, we encounter a no-go result indicating that such cooling cannot occur.
Finally, let us mention that there have been several studies on the problem of cooling, either using the resource-theoretic approach to thermodynamics or under a different umbrella of quantum thermodynamics [64, 65, 66, 67]. In particular, Ref. [22] uses the approach of catalytic thermal operations for the task of cooling. However, their focus is on providing sufficient and necessary conditions on the amount of resources needed to cool a system close to its ground state. Additionally, the catalyst is allowed to return $\epsilon$-away from its initial state. Therefore, the results presented in this section complement the previous ones by finding the optimal bounds on the heat exchange, which leads to cooling the initial system.
V Summary and Discussion
@@
In this paper, we explore the fundamental limits on state achievability with the aid of a strict catalyst under a thermal process. We propose an approach to address this question by constructing the sets of states that can be achieved from a given energy-incoherent state under catalytic thermal operations, as well as the sets of states that can, through catalytic thermal operations, achieve the initial state. These regions naturally highlight the advantages of using a strict catalyst in general thermodynamic processes, whose only assumption is energy conservation. Our construction enables us to characterise the catalysable past and future for main systems of arbitrary dimensions. This analytical approach is based solely on thermomajorisation relations and calculation of slopes of the resulting thermomajorisation curves, providing a way of verifying non-catalysability in finite number of steps. Furthermore, we established general bounds on the catalyst dimension for thermal operations and demonstrated that catalysis is effective under thermal operations even for systems of dimensionality as small as $d=3$. The catalytic advantages were quantified by the volume of the catalysable future, with a detailed discussion on its behaviour as a function of temperature. Moreover, we applied our findings to address two practical questions: the generation of entanglement under thermal operations and the optimal cooling of a quantum system. In the first application, we showed that thermal processes, which were once incapable of generating entanglement, allows its generation when a strict catalyst is employed. While in the second, we discussed the ultimate limits of cooling aided by catalysts.
@@
To address this problem, we overcome several challenges. Firstly, there is no efficient method for characterizing the set of allowed state transformations under catalytic thermal operations. A comprehensive analysis for inexact catalysis (where the catalyst is allowed to be returned with some minor error) was performed in [23]; while, in Ref. [28], the authors solved the characterization problem for three-dimensional systems and for a subset of initial states of generic dimensions in the context of catalysis in elementary thermal operations. Secondly, keeping track of the the set of achievable state after “decoupling” the catalyst without specifying the precise form of the catalyst state required by the transformation is a highly difficult mathematical problem. Approaching from the perspective of lattice theory, this problem can be framed by asking (i) how to characterise the higher-dimensional lattice for composite systems (main system plus catalyst) and (ii) how to project back to the native lattice of the main system while preserving the catalytic effects in the native lattice of the initial state. For small-dimensional systems (such as $d=3$), the entire convex polytope $\boldsymbol{p}\otimes\boldsymbol{c}$ can be constructed, and intersections between such a polytope and plane equations enforcing the catalyst recovery condition can be numerically identified. However, this method does not scale well and depends heavily on the initial state of the system and the catalyst. In contrast, our approach requires only constructing a set of quasi-probability vectors or determining the extremal points through thermomajorisation relations.
@@
There are many possibilities for extending our results. For instance, they can be leveraged to study catalysis within the framework of elementary thermal operations. Note that the characterization of the catalysable past and future relies entirely on the concept of tangent vectors, which tightly thermomajorise $\boldsymbol{p}$. On the other hand, the extreme points of set of states achievable under elementary thermal operations are known to be $\beta$-swaps that tightly thermomajorise the initial distribution. Consequently, one could explore how to adapt the technical aspects of our findings to this specific context. In a similar spirit, we could relax the condition requiring the catalyst to be returned unperturbed and allow for some small error. A natural quantifier for this error is the trace distance, and the problem could be reframed by asking for the set of states achievable with a given catalyst up to this specified amount of error [23].
It is tempting to suggest that the results obtained in this work, as an extension of prior results of Grabowecky and Gour [45], may point to even more general statements. Noting that majorisation and thermomajorisation curves are, in fact, totally ordered set of monotones, one may imagine the following statement: given two states, $X$ and $Y$, are such that the first and last monotones for $X$ are greater than that of $Y$, then the transformation is catalysable, thus providing a method for constructing the catalysable past and future for more general resource theories. This, however, is beyond the scope of this paper.
@@ Acknowledgements.
We thank Karol Życzkowski for asking a stimulating question at the beginning of the authors’ PhD, which led to this work, and for his fruitful comments on the first version of this manuscript. We also thank Kamil Korzekwa for valuable discussions that significantly contributed to the execution of this project, as well as for his insightful comments during the final stages of preparation of this manuscript. We acknowledge the financial support from Danish National Research Foundation’s bigQ grant (DNRF 142) is gratefully acknowledged. JCz acknowledges financial support by NCN PRELUDIUM BIS no. DEC-2019/35/O/ST2/01049 and “Research support module” as part of the “Excellence Initiative – Research University” program at the Jagiellonian University in Kraków. AOJ acknowledges financial support from VILLUM FONDEN through a research grant (40864) and EU Horizon Europe project QSNP (grant agreement no. 101114043).
@@ B.2 Embedding map approach
$\hat{\boldsymbol{p}}=\Biggl{[}\underbrace{\frac{p_{1}}{D_{1}},\ldots,\frac{p_{1}}{D_{1}}}_{D_{1}\>\text{times}},\underbrace{\frac{p_{2}}{D_{2}},\ldots,\frac{p_{2}}{D_{2}}}_{D_{2}\>\text{times}},\ldots,\underbrace{\frac{p_{d}}{D_{d}},\ldots,\frac{p_{d}}{D_{d}}}_{D_{d}\>\text{times}}\Biggr{]}.$ (71)